### v2

@@ Abstract
deep learning |||| ordinary differential equations |||| multiscale sampling |||| high dimension
@@ 1 Introduction
In this work, we propose a novel DeePODE method to establish global surrogate model for efficiently predicting multiscale dynamical systems. The DeePODE method takes on the ”local dynamical behavior” advantage through two components: the evolutionary Monte Carlo sampling method (EMCS) and the end-to-end deep neural network (DNN). The EMCS method combines the strengths of traditional Monte Carlo approaches with the system’s natural evolutionary dynamics. More specifically, the method uses Monte Carlo sampling to achieve global coverage of the high-dimensional space, addressing the curse of dimensionality. This global sampling is then enhanced by evolving each sample point along its local ODE trajectory at a speed determined by the system’s characteristic timescales. This hybrid approach naturally incorporates multiscale information into the sampling process. In regions where the system evolves slowly (low gradient/frequency regions), the method automatically collects fewer samples since large time steps are adopted to depict such domains. This hybrid sampling strategy efficiently generates representative data across the global high-dimensional space while respecting the system’s inherent behavior. Then, the DNN is designed to predict system dynamics using arbitrarily large time steps. This network embeds temporal evolution patterns of wide frequency spectrum, enabling accurate predictions over extended time intervals that would be computationally expensive with conventional numerical methods.
@@
We validate the effectiveness of DeePODE in various multiscale systems, including a predator-prey model, an electronic dynamical process, a battery electrolyte auto-ignition, and several turbulent reaction-diffusion systems considering detailed chemical kinetics. These systems represent critical areas of scientific and engineering interest with wide-ranging applications, in which DeePODE achieves satisfactory accuracy, robustness and computational efficiency. One particularly challenging area where DeePODE excels is in reaction-diffusion system (i.e. combustion) simulations [27]. The complexity of these simulations stems from multiple factors: the high dimensionality of chemical reaction processes involving numerous species, time scales spanning eight orders of magnitude, and the influence of turbulent flow fields. Such system is crucial for understanding coupled reaction and transport phenomenon, from propulsion system to atmospheric chemistry and interstellar reactions. DeePODE demonstrates remarkable capabilities in addressing these challenges. One of its key strengths lies in the global surrogate it provides for combustion simulation—a single pretrained ODE surrogate model that can seamlessly couple with 1D, 2D, or 3D PDEs under previously unseen initial conditions, without requiring retraining or fine-tuning. Another advantage is the adaptability of the model. DeePODE models can be readily integrated into different numerical platforms, offering significant computational speed improvements—often two orders of magnitude faster than traditional solvers. This combination of accuracy, acceleration, and adaptability makes DeePODE a powerful tool for complex system simulations, such as biomass conversions [28] and air pollution predictions [29].
@@ 2 Methodology
Figure 1: Schematic diagram of the DeePODE method. (A) Flowchart of DeePODE method. After obtaining a manifold dataset and an MC dataset, each point of the MC dataset is taken as the initial condition of a corresponding dynamical system. Sample along the evolution trajectory of MC data sparsely and filter them by the temporal gradient range of the manifold dataset. The remaining data composes the final EMCS dataset. A DNN is used to fit the dataset. Our DeePODE method can be regarded as EMCS combined with DNN, namely ”DeePODE = EMCS + DNN” as in the annotation. (B) Example of prey-predator system. The left and right figures represent the data distribution obtained by the Monte-Carlo and the EMCS methods, respectively, and the color of the data points represents the change’s norm. The gray regions in the second and third figures represent some evolution trajectories of this system. The yellow points mean the sampled point in the EMCS method. (C) Example of Chemical kinetic system. Three figures serve the same purpose as the corresponding figures in (B) except that $x_{1}$ and $x_{2}$ are the data projected onto the first two principal directions of the manifold dataset.
@@ 2.2 Evolutionary Monte Carlo sampling
Due to the unique nature of each ODE system, it is important to understand its characteristic data distribution at the beginning. The range estimation process addresses this need by determining the system’s feasible ranges through long-term evolution from several extreme initial conditions of interest, as shown in Fig. 1A(a). The dataset structure can be visualized through paired yellow dots and green columns. Each yellow dot represents a phase state $\boldsymbol{x}(t)$ in the high-dimensional state space, while its corresponding green column represents the temporal change $\boldsymbol{u}(t)$ at that state. Together, they form a training pair $(\boldsymbol{x}(t),\boldsymbol{u}(t))$, where $\boldsymbol{x}(t)$ serves as the input datapoint and $\boldsymbol{u}(t)$ as its label. By tracking these trajectories and collecting samples along them, we obtain a rough but essential estimation of the feasible range $[x^{\min}_{i},x^{\max}_{i}]$ and $[u^{\min}_{i},u^{\max}_{i}]$ for each component $x_{i}$ in the state vector $\boldsymbol{x}_{t}$. This initial dataset, denoted as $\mathbb{D}_{\text{Init}}:={(\boldsymbol{x}(t),\boldsymbol{u}(t))}$, provides crucial information about the system’s bounds. A typical complex chemical reaction system like electrolyte thermal runaway would involve simulating electrolytes from various extreme temperature and concentration combinations to understand the full range of possible data distribution. This comprehensive range estimation is essential because electrolyte thermal runaway involves complex interactions between temperature-dependent decomposition reactions, heat generation, and species evolution, each operating across different scales and potentially leading to critical transitions in system behavior.
@@
The final step, evolution augmented generation, enhances the dataset by evolving each Monte Carlo sample point along its natural trajectory according to the ODEs as shown in Fig. 1A(c). By sampling data points from these local trajectories, we effectively capture the system’s behavior in the manifold surrounding each Monte Carlo point. A critical insight in this evolution step is the use of non-uniform evolution time steps $\Delta t_{i}=\boldsymbol{\tau_{i}}=[\tau_{1},\tau_{2},\cdots,\tau_{k}]$, which enables the updated dataset $\mathbb{D}_{\rm EMCS}$ to capture more characteristic timescales of the system. By allowing the sampling intervals to vary based on local dynamics, we can efficiently represent both rapid changes and slow evolution within the system.
@@ 3.1 Two-dimensional predator-prey model
The performance comparison reveals distinct outcomes across different methods. As shown in Fig. 2(A), the DNNs trained using the MC sampling or manifold sampling show substantial deviations from the reference solution. In contrast, DeePODE maintains accurate predictions even over extended time periods (up to t=100 seconds). Moreover, DeePODE achieves this accuracy while using larger integration steps than the explicit Euler method, demonstrating a crucial advantage: it combines the computational efficiency of explicit schemes with enhanced numerical stability. This combination of accuracy and efficiency makes DeePODE particularly valuable for long-time integration of dynamical systems.
Figure 2: DeePODE for multiscale dynamic systems. (A) Two-dimensional predator-prey model. DNN prediction and direct integration results for $x_{1}$ from the initial value $x_{1}=3$, $x_{2}=2$. RK45, MC, MF, and EE represent Runge-Kutta-Fehlberg scheme, Monte Carlo, manifold sampling methods, and Explicit Euler scheme, respectively. (B) Ring modulator model. DNN prediction and direct integration results for $U_{1}$ and $I_{2}$. (C) Battery electrolyte temporal evolution of temperature and radical HCO mass fraction. The initial condition is $p_{0}=1$ atm, $T_{0}=1400$ K, equivalence ratio $\phi_{0}=1$.
@@ 3.3 Battery electrolyte thermal runaway
[…]
@@ 3.4 Reaction-diffusion system with complex chemical kinetics
In this section, we evaluate DeePODE models using several representative chemical kinetic models, including methane oxidation model GRI-Mech 3.0, which comprises 53 species and 325 reactions; DRM19, with 21 species and 84 reactions for methane/air reaction; and an n-heptane oxidation model containing 34 species and 191 reactions [34]. We train a corresponding DeePODE model for each mechanism, which we have made publicly available through our repository (https://github.com/intelligent-algorithm-team/intelligent-combustion). These models are then thoroughly evaluated using various flow benchmark tests. A key strength of DeePODE is its versatility - once trained, a single model can be directly integrated into one-, two-, or three-dimensional flow simulations without requiring additional training or adjustment. This ”train once, use anywhere” capability significantly enhances its practical reactor applications. The CFD code used for algorithm tests is EBI-DNS [35]. To demonstrate the portability of DeePODE, we integrate the DNN model into two widely-used CFD frameworks: ASURF [36, 37] and DeepFlame [38] codes. Further details on chemical kinetics sampling and additional examples are provided in the Appendix B.2.
@@ 3.5.2 Computational efficiency
The computational performance data in Table 1 demonstrates DeePODE’s superior efficiency across multiple test cases. The method achieves speedup ratios of 10100×10-100$\times$10 - 100 × compared to traditional direct integration methods, showing impressive gains in both point-wise and parallel computing configurations. The most significant acceleration is observed in reaction-diffusion systems, where DeePODE exhibits its greatest performance advantages. These quantitative results validate DeePODE’s ability to substantially improve computational efficiency while preserving solution quality, making it a powerful alternative to conventional approaches.
@@ 4 Discussion
The superior performance of EMCS in addressing multiscale dynamical systems arises from its innovative approach to adaptive data sampling and temporal scale coverage. Through its evolutionary process, EMCS adapts to the local gradient structure of the phase space, efficiently allocating samples where they are most needed. This approach generates fewer data points in regions of low gradient—where system behavior is more predictable—aligning with neural networks’ natural ability to learn smooth, low-frequency patterns. As a result, DeePODE achieves accurate approximations across diverse operating conditions while maintaining an economical dataset size. Another distinctive strength of EMCS is its comprehensive coverage of characteristic timescales in high-dimensional systems, leading to robust generalization in temporal evolution prediction. As evidenced in Fig. 6(D), EMCS samples effectively across the full spectrum of system dynamics, whereas traditional Monte Carlo methods fail to capture this temporal diversity. This thorough representation of multiple timescales fundamentally enhances the method’s predictive accuracy and reliability.
@@ Acknowledgements
This work is sponsored by the National Key R&D Program of China Grant No. 2022YFA1008200 (Z. X., Y. Z.), the National Natural Science Foundation of China 92470127(T. Z.), No. 62002221 (Z. X.), 12371511 (ZX), 12422119 (ZX), 92270203 (T. Z.), Grant No. 12101402 (Y. Z.), the Lingang Laboratory Grant No.LG-QS-202202-08 (Y. Z.), Shanghai Municipal of Science and Technology Project Grant No. 20JC1419500 (Y. Z.), Shanghai Municipal of Science and Technology Major Project No. 2021SHZDZX0102, and the HPC of School of Mathematical Sciences and the Student Innovation Center, and the Siyuan-1 cluster supported by the Center for High-Performance Computing at Shanghai Jiao Tong University, Key Laboratory of Marine Intelligent Equipment and System, Ministry of Education, P.R. China. SJTU Kunpeng&Ascend Center of Excellence partially supported this work.
Competing interest
The authors declare no competing interests.
Data availability
All data supporting the findings of this study will be available upon publication.
Code availability
The source code will be available at https://github.com/intelligent-algorithm-team/intelligent-combustion upon publication.
@@ A.1 EMCS Method
0:  Initial conditions $\boldsymbol{x}_{0}$, evolution time distribution $\mathcal{T}$ and evolution steps $k$;
1:  $X_{\rm MF}=\{\boldsymbol{x}(i\cdot\delta t)\}_{i}$ where $\boldsymbol{x}(0)=\boldsymbol{x}_{0}$;
2:  $R_{\rm MF}=[\min X_{\rm MF},\max X_{\rm MF}]\times[\min u_{\rm MF},\max u_{\rm MF}]$;
3:  $\mathcal{F}_{\rm MF}=\mathcal{U}(\min X_{\rm MF},\max X_{\rm MF})$;
4:  $X_{\rm EMCS,0}\sim\mathcal{F}_{\rm MF}$;
5:  for $i=1,2,\cdots,k$ do
6:     $\tau_{i}\sim\mathcal{T}$;
[…]
8:     $\mathbb{D}_{\rm EMCS,i}=\left(X_{\rm EMCS,i},u_{\rm EMCS,i}\right)\cap R_{\rm MF}$
9:  end for
10:  return $\mathbb{D}_{\rm EMCS}=\cup_{i=0}^{n}\mathbb{D}_{\rm EMCS,i}$.

### v3

@@ Abstract
deep learning , ordinary differential equations , multiscale sampling , high dimension
journal: Computer Physics Communications
[…]
[4]organization=AI for Science Institute,city=Beijing,country=China
[…]
[7]organization=Key Laboratory of Spacecraft Design Optimization and Dynamic Simulation Technologies, Ministry of Education,postcode=102206, city=Beijing, country=China
[…]
[8]organization=Key Laboratory of Marine Intelligent Equipment and System, Ministry of Education,postcode=201100,country=China
@@ 1 Introduction
In this work, we propose a novel DeePODE method to establish global surrogate model for efficiently predicting multiscale dynamical systems. The DeePODE method takes on the ”local dynamical behavior” advantage through two components: the evolutionary Monte Carlo sampling method (EMCS) and the end-to-end deep neural network (DNN). The EMCS method combines the strengths of traditional Monte Carlo approaches with the system’s natural evolutionary dynamics. More specifically, the method uses Monte Carlo sampling to achieve global coverage of the high-dimensional space, addressing the curse of dimensionality. This global sampling is then enhanced by evolving each sample point along its local ODE trajectory at a speed determined by the system’s characteristic timescales. This hybrid approach naturally incorporates multiscale information into the sampling process. In regions where the system evolves slowly (low gradient regions), the method automatically collects fewer samples since large time steps are adopted to depict such domains. This hybrid sampling strategy efficiently generates representative data across the global high-dimensional space while respecting the system’s inherent behavior. Then, the DNN is designed to predict system dynamics using arbitrarily reasonable large time steps. This network embeds temporal evolution patterns of wide frequency spectrum, enabling accurate predictions over extended time intervals that would be computationally expensive with conventional numerical methods. This work focuses on demonstrating the importance of sampling representative data, which enable neural networks, even a vanilla fully-connected network, to achieve high accuracy in simulating high-dimensional and stiff ODEs.
@@
We validate the effectiveness of DeePODE in various multiscale systems, including a predator-prey model, an electronic dynamical process, a battery electrolyte auto-ignition, and several turbulent reaction-diffusion systems considering detailed chemical kinetics. These systems represent critical areas of scientific and engineering interest with wide-ranging applications, in which DeePODE achieves satisfactory accuracy, robustness and computational efficiency. One particularly challenging area where DeePODE excels is in reaction-diffusion system (i.e. combustion) simulations [27]. The complexity of these simulations stems from multiple factors: the high dimensionality of chemical reaction processes involving numerous species, time scales spanning eight orders of magnitude, and the influence of turbulent flow fields. Such system is crucial for understanding coupled reaction and transport phenomenon, from propulsion system to atmospheric chemistry and interstellar reactions.
DeePODE demonstrates remarkable capabilities in addressing these challenges. One of its key strengths lies in the global surrogate it provides for combustion simulation—a single pretrained ODE surrogate model that can seamlessly couple with 1D, 2D, or 3D PDEs under previously unseen initial conditions, without requiring retraining or fine-tuning. Another advantage is the adaptability of the model. DeePODE models can be readily integrated into different numerical platforms, offering significant computational speed improvements—often two orders of magnitude faster than traditional solvers. This combination of accuracy, acceleration, and adaptability makes DeePODE a powerful tool for complex system simulations, such as biomass conversions [28] and air pollution predictions [29].
@@ 2 Methodology
Figure 1: Schematic diagram of the DeePODE method. (A) Flowchart of DeePODE method. After obtaining a manifold dataset and an MC dataset, each point of the MC dataset is taken as the initial condition of a corresponding dynamical system. Sample along the evolution trajectory of MC data sparsely and filter them by the temporal gradient range of the manifold dataset. The remaining data composes the final EMCS dataset. A DNN is used to fit the dataset. Our DeePODE method can be regarded as EMCS combined with DNN, namely ”DeePODE = EMCS + DNN” as in the annotation. (B) Example of prey-predator system. The left and right figures represent the data distribution obtained by the Monte-Carlo and the EMCS methods, respectively, and the color of the data points represents the change’s norm. The smooth gray regions in the second and third figures represent some evolution trajectories of this system. The yellow circular markers denote the sampled points in the EMCS method. (C) Example of Chemical kinetic system. Three figures serve the same purpose as the corresponding figures in (B) except that $x_{1}$ and $x_{2}$ are the data projected onto the first two principal directions of the manifold dataset. (For interpretation of the colors in the figure(s), the reader is referred to the web version of this article.)
@@ 2.2 Evolutionary Monte Carlo sampling
Due to the unique nature of each ODE system, it is important to understand its characteristic data distribution at the beginning. The range estimation process addresses this need by determining the system’s feasible ranges through long-term evolution from several extreme initial conditions of interest. To account for the diverse scales of dynamical systems (e.g., species concentrations in chemical kinetics), we randomly select physically realistic initial conditions based on the system’s physical constraints—for example, in chemical reaction kinetics, only fuel and oxidizer are present initially, with other species’ mass fractions set to zero, and temperatures are capped at realistic values (e.g., 3000 K). These conditions are evolved numerically using CVODE with a small time step (e.g., $10^{-8}$ s) until stabilization or periodic behavior emerges. The resulting evolutionary trajectories form the dataset $\mathbb{D}_{\text{init}}=\{\langle\boldsymbol{x}(t),\boldsymbol{u}(t)\rangle\}$, where $\boldsymbol{u}(t):=\boldsymbol{x}(t+\Delta t)-\boldsymbol{x}(t)$ is the temporal change. This process, requiring only  1% of the full training data volume, provides an efficient and robust range estimation for subsequent sampling steps. As shown in Fig. 1A(a), the dataset structure can be visualized through paired yellow dots and green columns. Each yellow dot represents a phase state $\boldsymbol{x}(t)$ in the high-dimensional state space, while its corresponding green column represents the temporal change $\boldsymbol{u}(t):=\boldsymbol{x}(t+\Delta t)-\boldsymbol{x}(t)$, where $\Delta t$ is the time step used for DNN prediction. Together, they form a training pair $(\boldsymbol{x}(t),\boldsymbol{u}(t))$, where $\boldsymbol{x}(t)$ serves as the input data point and $\boldsymbol{u}(t)$ as its label. By tracking these trajectories and collecting samples along them, we obtain a rough but essential estimation of the feasible range $x^{\min}_{i},x^{\max}_{i}]$ and $u^{\min}_{i},u^{\max}_{i}]$ for each component $x_{i}$ in the state vector $\boldsymbol{x}_{t}$. This initial dataset, denoted as $\mathbb{D}_{\text{Init}}:={(\boldsymbol{x}(t),\boldsymbol{u}(t))}$, provides crucial information about the system’s bounds. A typical complex chemical reaction system like electrolyte thermal runaway would involve simulating electrolytes from various extreme temperature and concentration combinations to understand the full range of possible data distribution. This comprehensive range estimation is essential because electrolyte thermal runaway involves complex interactions between temperature-dependent decomposition reactions, heat generation, and species evolution, each operating across different scales and potentially leading to critical transitions in system behavior.
@@
The final step, evolution augmented generation, enhances the dataset by evolving each Monte Carlo sample point along its natural trajectory according to the ODEs as shown in Fig. 1A(c). By sampling data points from these local trajectories, we effectively capture the system’s behavior in the manifold surrounding each Monte Carlo point. A critical insight in this evolution step is the use of non-uniform evolution time steps $\Delta t_{i}=\boldsymbol{\tau_{i}}=[\tau_{1},\tau_{2},\cdots,\tau_{k}]$, which enables the updated dataset $\mathbb{D}_{\rm EMCS}$ to capture more characteristic timescales of the system. By allowing the sampling intervals to vary based on local dynamics, we can efficiently represent both rapid changes and slow evolution within the system. In practice, we employ an increasing $\tau_{i}$ strategy where $\tau_{i+1}\geq\tau_{i}$. The selection criteria for $\tau_{i}$ are discussed in detail in Section 3.5.3.
@@ 3.1 Two-dimensional predator-prey model
The performance comparison reveals distinct outcomes across different methods. As shown in Fig. 2(A), the DNNs trained using the MC sampling or manifold sampling show substantial deviations from the reference solution. In contrast, DeePODE maintains accurate predictions even over extended time periods (up to t=100 seconds). Although all three sampling methods use datasets of identical size, they produce distinct data distributions (Fig. 1(B)). This demonstrates that the spatial distribution of training data critically influences DNN performance, with EMCS achieving broader and more representative coverage of the phase space. Moreover, DeePODE achieves this accuracy while using larger integration steps than the explicit Euler method, demonstrating a crucial advantage: it combines the computational efficiency of explicit schemes with enhanced numerical stability. This combination of accuracy and efficiency makes DeePODE particularly valuable for long-time integration of dynamical systems.
Figure 2: DeePODE for multiscale dynamic systems. (A) Two-dimensional predator-prey model. DNN prediction and direct integration results for $x_{1}$ from the initial value $x_{1}=3$, $x_{2}=2$. RK45, MC, MF, and $E$ represent Runge-Kutta-Fehlberg scheme, Monte Carlo, manifold sampling methods, and Explicit Euler scheme, respectively. (B) Ring modulator model. DNN prediction and direct integration results for $U_{1}$ and $I_{2}$. (C) Battery electrolyte temporal evolution of temperature and radical HCO mass fraction. The initial condition is $p_{0}=1$ atm, $T_{0}=1400$ K, equivalence ratio $\phi_{0}=1$.
@@ 3.3 Battery electrolyte thermal runaway
We utilize Cantera to obtain the reference solution and generate datasets for DNN models. More specifically, for the MC method, 500,000 data points are sampled uniformly within the range $\text{temperature}\in[300,3000]$ K, $\text{pressure}\in[0.5,3]$ atm, $\text{species mass fraction}\in[0,1]$. For the DeePODE method, we randomly pick 100,000 data points within the same range and apply EMCS for those points with evolution time sequence $\boldsymbol{\tau}=[10^{-6},10^{-5},10^{-4},10^{-3}]$ s. The corresponding label datasets are obtained by evolving the input vectors for $\Delta t=10^{-6}$ s. We use the two datasets to train the corresponding DNNs with 1600,800,400 hidden neurons for three hidden layers, respectively. With the stiffness constraint, the integration time step inside Cantera is $10^{-8}$, while the DNN models predict the state after $\Delta t=10^{-6}$ given an input state.
@@ 3.4 Reaction-diffusion system with complex chemical kinetics
In this section, we evaluate DeePODE models using several representative chemical kinetic models, including methane oxidation model GRI-Mech 3.0, which comprises 53 species and 325 reactions; DRM19, with 21 species and 84 reactions for methane/air reaction; and an n-heptane oxidation model containing 34 species and 191 reactions [34]. We train a corresponding DeePODE model for each mechanism, which we have made publicly available through our repository (https://github.com/intelligent-algorithm-team/intelligent-combustion). These models are then thoroughly evaluated using various flow benchmark tests. A key strength of DeePODE is its versatility - once trained, a single model can be directly integrated into one-, two-, or three-dimensional flow simulations without requiring additional training or adjustment. This ”train once, use anywhere” capability significantly enhances its practical reactor applications. The CFD code used for algorithm tests is EBI-DNS [35]. To demonstrate the portability of DeePODE, we integrate the DNN model into two widely-used CFD frameworks: ASURF [36, 37] and DeepFlame [38] codes. The ODE solver used for both generating training data and evaluation is CVODE[39], which employs the implicit Backward Differentiation Formula (BDF) method, designed for robust handling of stiff systems. Further details on chemical kinetics sampling and additional examples are provided in the Appendix B.2.
@@ 3.5.2 Computational efficiency
The computational performance data in Table 1 demonstrates DeePODE’s superior efficiency across multiple test cases. The method achieves speedup ratios of $10-100\times$ compared to traditional direct integration methods, showing impressive gains in both point-wise and parallel computing configurations. The most significant acceleration is observed in reaction-diffusion systems, where DeePODE exhibits its greatest performance advantages. These quantitative results validate DeePODE’s ability to substantially improve computational efficiency while preserving solution quality, making it a powerful alternative to conventional approaches.
@@ 4 Discussion
The superior performance of EMCS in addressing multiscale dynamical systems arises from its innovative approach to adaptive data sampling and temporal scale coverage. Through its evolutionary process, EMCS adapts to the local gradient structure of the phase space, efficiently allocating samples where they are most needed. This approach generates fewer data points in regions of low gradient—where system behavior is more predictable—aligning with neural networks’ natural ability to learn smooth, low-frequency patterns. As a result, DeePODE achieves accurate approximations across diverse operating conditions while maintaining an economical dataset size. Another distinctive strength of EMCS is its comprehensive coverage of characteristic timescales in high-dimensional systems, leading to robust generalization in temporal evolution prediction. As evidenced in Fig. 6(D), EMCS samples effectively across the full spectrum of system dynamics, whereas traditional Monte Carlo methods fail to capture this temporal diversity. This thorough representation of multiple timescales fundamentally enhances the method’s predictive accuracy and reliability. It’s noted that in deep learning methodologies, dataset size typically has a significant influence on model performance. Under appropriate conditions, larger training datasets generally yield improved model outcomes – a principle that equally applies to the EMCS sampling approach.
@@
Declaration of interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI in scientific writing
During the preparation of this work the authors used generative AI in order to polish the language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Acknowledgements
This work is sponsored by the National Natural Science Foundation of China 92470127 (T. Z.), No. 62002221 (Z. X.), 12371511 (Z. X.), 12422119 (Z. X.), 92270203 (T. Z.), Grant No. 12101402 (Y. Z.), the Lingang Laboratory Grant No.LG-QS-202202-08 (Y. Z.), Shanghai Municipal of Science and Technology Project Grant No. 20JC1419500 (Y. Z.), Shanghai Municipal of Science and Technology Major Project No. 2021SHZDZX0102, and the HPC of School of Mathematical Sciences and the Student Innovation Center, and the Siyuan-1 cluster supported by the Center for High-Performance Computing at Shanghai Jiao Tong University, Key Laboratory of Marine Intelligent Equipment and System, Ministry of Education, P.R. China. SJTU Kunpeng & Ascend Center of Excellence partially supported this work.
@@ A.1 EMCS Method
0: System dimension $n$, the data size $N_{0}$ of MC sampling, the evolution time selection strategy $\boldsymbol{\tau}=[\tau_{1},\tau_{2},\cdots,\tau_{k}]$, the time step for DNN prediction $\Delta t$.
1: // Step 1: Range Estimation
2: Collect data points from multiple temporal trajectories and form the set $\boldsymbol{x}_{\rm MF}$;
3: Compute the label set $\boldsymbol{u}_{\rm MF}$;
4: Update the range estimation $R_{\rm MF}=[\min\boldsymbol{x}_{\rm MF},\max\boldsymbol{x}_{\rm MF}]\times[\min\boldsymbol{u}_{\rm MF},\max\boldsymbol{u}_{\rm MF}]$;
5: // Step 2: Monte Carlo (MC) Sampling
6:for $i=1:n$ do
7:  if  the component $x^{i}$ spans multiple scales  then
8:   Sample $N_{0}$ sub-dimension samples $\log x^{i}\sim\mathcal{U}(\log\min\boldsymbol{x}^{i}_{\rm MF},\log\max\boldsymbol{x}^{i}_{\rm MF})$;
9:  else
10:   Sample $N_{0}$ sub-dimension samples $x^{i}\sim\mathcal{U}(\min\boldsymbol{x}^{i}_{\rm MF},\max\boldsymbol{x}^{i}_{\rm MF})$;
11:  end if
12:end for
13: Merge all the $N_{0}$ data samples into the set $\mathbb{D}_{\text{MC}}$.
14: // Step 3: Evolution Augmented Generation
15:for  each $\boldsymbol{x}(t)\in\mathbb{D}_{\text{MC}}$ do
16:  $\boldsymbol{x}_{1}:=\boldsymbol{x}(t)$;
17:  for $i=1:k$ do
18:   $\boldsymbol{x}_{i+1}:=\boldsymbol{x}_{i}+\int_{t}^{t+\tau_{i}}\boldsymbol{f}(\boldsymbol{x},s)\mathrm{d}s$. $\quad\triangleright~\text{temporal evolution}$;
19:  end for
20:  for $i=1:k+1$ do
21:   $\boldsymbol{u}_{i}:=\int_{t}^{t+\Delta t}\boldsymbol{f}(\boldsymbol{x},s)\mathrm{d}s$; $\quad\triangleright~\text{label generation}$
22:   Update the dataset $\mathbb{D}_{\text{EMCS}}=\mathbb{D}_{\text{EMCS}}\cup\{(\boldsymbol{x}_{i},\boldsymbol{u}_{i})\}\cap R_{\rm MF}$.
23:  end for
24:end for
24: The final dataset $\mathbb{D}_{\text{EMCS}}$.